Station Messages
18 Jul 2018 02:53 UTC 2018199+02:53 UTC

Overview Datums Bench Marks Data Query Site Visit Daily QC Messages Attributes

127: GBRA Station #1

28° 15' 35" N   96° 46' 25" W
GBRA1
Guadalupe Blanco River Authority (GBRA) GBRA

2012200+0601 (info, ndurham)
trb : now (07/18/2018 02:53 - 07/18/2018 02:54) :

For the following date range, 2004054 to 2012193, Hydrolab turbidity data was found to be erroneous. The corrected data was reimported on July 17, 2012.

2015352+1445 (note, hmahlke)
ALL :

low data recovery

2018198+1749 (retired, llloyd)
dth : 2018178+1600-2018193+1654 (06/27/2018 16:00 - 07/12/2018 16:54) :

Datasonde deployed from 2018178+1600 to 2018193+1654 was not equipped with a depth probe.

2018158+1713 (retired, dburch)
trb : 2018142+0506-2018156+0330 (05/22/2018 05:06 - 06/05/2018 03:30) :

Turbidity values higher than expected from 2018142+0506-2018156+0330. High values expected to be from biofouling.

2018158+1658 (retired, dburch)
sal,spcon,sat,do : 2018115+2282-2018125+1788 (04/25/2018 23:22 - 05/05/2018 18:28) :

Data values lower than expected from 2018115+2282-2018125+1788. Biofouling expected cause.

2018073+2009 (retired, lloyd)
trb : 2018055+1212-2018067+1518 (02/24/2018 12:12 - 03/08/2018 15:18) :

Higher than expected turbidity values starting 2018055+1212. The datasonde was replaced on 2018067+1518. Biofouling was the suspected cause. Suspect data was removed.

2018053+1932 (retired, lloyd)
ALL : 2018051+2030 (02/20/2018 20:30 - 02/20/2018 20:31) :

No data dissemination starting 2018051+2030. IP modem not checking in. Connected to station modem manually at 2018053+0025 and issued a reset. Data dissemination resumed and logged data from the station were automatically disseminated as of 2018053+1926.

2017318+1632 (retired, lloyd)
dth : 2017286+1424 (10/13/2017 14:24 - 10/13/2017 14:25) :

Backup datasonde installed on 2017286+1424 is not equipped with a depth sensor

2017286+1741 (retired, lloyd)
ALL : 2017280+1154 (10/07/2017 11:54 - 10/07/2017 11:55) :

No data dissemination starting 2017280+1154, cannot establish a connection to the cellular modem

2017282+2026 (retired, lloyd)
vlz,ssx,ssz,sdth,vlx,vly,ssy,sbat : 2016322+1636 (11/17/2016 16:36 - 11/17/2016 16:37) :

The current meter is not responding to commands and is in the process of being evaluated/repaired by the manufacturer

2016347+1446 (retired, lloyd)
ALL : 2016344+0154 (12/09/2016 01:54 - 12/09/2016 01:55) :

No data being disseminated starting 2016344+0154

2016340+2155 (retired, lloyd)
vlz,ssx,ssz,sdth,vlx,vly : Little data dissemination starting 2016270+0712 ( - 01/01/1970 00:00) :

Little data dissemination from current meter starting 2016270+0712

2016340+2154 (retired, lloyd)
sdth : 2016287+0536 (10/13/2016 05:36 - 10/13/2016 05:37) :

Current meter depth data suspect starting 2016287+0536, flatlining on -0.001

2016249+1701 (retired, lloyd)
ssx,sdth,vlx,vly,ssy,sbat : 2016245+0330 (09/01/2016 03:30 - 09/01/2016 03:31) :

No current meter data dissemination starting 2016245+0330

2016218+1209 (retired, lloyd)
ph,do,spcon,dth,bat,sal,trb,sat : 2016216+0212 (08/03/2016 02:12 - 08/03/2016 02:13) :

No data disseminated from the datasonde starting 2016216+0212

2016153+1958 (retired, lloyd)

No datasonde data being disseminated starting 2016152+2312.

2016087+1039 (retired, jrizzo)
ssx,vlx,vly,ssy : 2016026+2218 (01/26/2016 22:18 - 01/26/2016 22:19) :

No current meter data starting 2016026+2218. Station reinstall pending

2016076+1214 (retired, jrizzo)
ALL : 2016074+0848 (03/14/2016 08:48 - 03/14/2016 08:49) :

No data being disseminated starting 2016074+0848

2016025+1608 (retired, jrizzo)

No current meter data

2016025+1608 (retired, jrizzo)
sbat : 2015362+2256 (12/28/2015 22:56 - 12/28/2015 22:57) :

Station reinstall pending Jan 2016

2016024+1334 (retired, jrizzo)

no current data

2015362+2256 (retired, lloyd)

Current meter data resumed, water quality data missing

2015356+1418 (retired, jrizzo)

No data

2015351+1529 (retired, lloyd)

No data

2015318+1401 (retired, jrizzo)

DO is zero

2015316+1422 (retired, blorentson)
2015308+1400 (11/04/2015 14:00 - 11/04/2015 14:01) :

No data dissemination starting 2015308+1400, power to building housing datalogger is turned off for maintenance. Power is scheduled to be returned on 11/9/15.

2015313+0344 (retired, lloyd)

No data

2015303+1612 (retired, lloyd)

Turb flatline high

2015295+1355 (retired, blorentson)

Turbidity flatline high @3000

2015289+0839 (retired, jrizzo)
ALL :

low data recovery

2015286+0218 (retired, lloyd)

Sal, con, turb @ 0, batt at 10.5v

2015283+1213 (retired, jrizzo)

DO flatlined at 0

2015281+1307 (retired, blorentson)

Turbidity flatline high @3000

2015278+0915 (retired, jrizzo)

Current meter data missing

2015266+1315 (retired, lloyd)

Salinity data reading "0".

2015255+1009 (retired, jrizzo)
trb :

TRB suspect, higher than expected values, flatlining near 300 nTU

2015245+1259 (retired, jrizzo)
sal,spcon : 2015236+2018 (08/24/2015 20:18 - 08/24/2015 20:19) :

Specific conductance/salinity suspect starting 2015236+2018. Probe malfunction suspect, waiting on replacement sensor.

2015238+1846 (retired, lloyd)
do,sat : 2015216+1630 (08/04/2015 16:30 - 08/04/2015 16:31) :

Sat and DO disabled starting 2015216+1630, probe malfunction suspect

2015238+1839 (retired, lloyd)

Data intermittent

2015217+1447 (retired, lloyd)
do,sat : 2015205+1600 (07/24/2015 16:00 - 07/24/2015 16:01) :

SAT and DO suspect starting 2015205+1600, lower than expected values

2015214+2259 (retired, lloyd)

No data

2015211+2031 (retired, blorentson)

No data

2015166+1301 (retired, blorentson)

No pH

2015155+1643 (retired, blorentson)

pH not reporting

2015148+1241 (retired, lloyd)
bat,sal,trb,ph,wtp,do,spcon,dth,sat : 2015141+1454 (05/21/2015 14:54 - 05/21/2015 14:55) :

No water quality data disseminated starting 2015141+1454

2015131+2038 (retired, lloyd)

Little data recovery

2015111+1937 (retired, lloyd)

No data

2015066+1311 (retired, jrizzo)

No data

2015039+0809 (retired, lloyd)

Battery 1 voltage dropped @ 6v

2015014+1447 (retired, lloyd)
sal,spcon : 2014360+2118 (12/26/2014 21:18 - 12/26/2014 21:19) :

spcon and sal suspect, lower than expected values, near 0

2015006+1152 (retired, amendoza)

Batt 1 @5V

2014339+2037 (retired, lloyd)

No data

2014337+1202 (retired, amendoza)

No current data

2014323+1157 (retired, amendoza)

No data dissemination due to drop in battery voltage.

2014321+0206 (retired, lloyd)

Intermittent data

2014258+0454 (retired, lloyd)

data intermittent

2014211+1154 (retired, amendoza)

No data

2014184+1314 (retired, lloyd)
2014166+1327 (retired, jrizzo)
ph :

no data

2014120+1109 (retired, amendoza)

No data

2014070+1140 (retired, amendoza)

Missing data from multiple series

2014050+1220 (retired, amendoza)

No data

2014047+0145 (retired, lloyd)
do,spcon,bat,sal,wtp,sat :

No data from multiple sensors

2014042+0556 (retired, lloyd)

No data, batteries at 5v

2014042+0556 (retired, lloyd)

No data

2014013+1410 (retired, amendoza)

Battery dead, no data

2013354+1451 (retired, lloyd)
ph : 12.03.2013 (12/03/2013 00:00 - 12/04/2013 00:00) :

pH disabled on datasonde installed on 12/3/13 due to probe malfunction

2013352+1633 (retired, lloyd)

No current meter data

2013352+1633 (retired, lloyd)
dth : 2013337+1736 (12/03/2013 17:36 - 12/03/2013 17:37) :

Water depth sensor suspect on datasonde installed on 2013337+1736

2013339+1709 (retired, lloyd)

No data

2013339+1708 (retired, lloyd)

Sporadic data

2013311+1154 (retired, amendoza)

Data sporadic

2013274+1037 (retired, amendoza)

No data being received due to datalogger malfunction. Replacement datalogger is being built

2013226+1116 (retired, amendoza)

No water depth data

2013162+0455 (retired, lloyd)

No data

2013154+1625 (retired, lloyd)

No data since 2013153+1700

2013088+0909 (retired, ussery)
ph,do,spcon,dth,sal,trb,sat : 2013085+1700 (03/26/2013 17:00 - 03/26/2013 17:01) :

Datasonde firmware malfunction on 26MAR13. Data has resumed.

2013086+0831 (retired, ussery)

No data from hydrolab since 2013085+1718.

2013014+0916 (retired, ussery)
ALL : 2013004+1530,now (01/04/2013 15:30 - 07/18/2018 02:53) :

No data since 2013004+1530 GMT. I tried to connect manually and didn't get a response from either the current meter or the YSI.

2012353+1442 (retired, lloyd)

Trubidity reading 0

2012263+1921 (retired, lloyd)

Battery voltage dropped. Battery was replaced on 2012263+1900

2012244+1010 (retired, ussery)

Data from current meter flatlining starting 2012240+0030. Current meter switched with working one.

2012240+1305 (retired, lloyd)

Sporadic data being received.

2012240+1303 (retired, lloyd)

Data from sontek flatlining.

2012226+0955 (retired, ussery)

No data from sontek since 8/4/2012.

2012219+1249 (retired, lloyd)

Turbidity data suspect starting 6/26/12

2012146+1026 (retired, ussery)

Turbidity readings flat-lining starting 5/2/2012. Replacement datasonde is being shipped.

2012117+1445 (retired, lloyd)

No data received from Sontek.

2012099+0146 (retired, lloyd)

No current meter data since 4/6/2012.

2012097+0605 (retired, lloyd)

No data from water quality instrument since 3/29/2012.

2012087+1336 (retired, lloyd)

No data from current meter.

2012087+1336 (retired, lloyd)

No data since 3/23/2012.

2012060+0231 (retired, lloyd)
ALL : 2012055+1400-2012058+1536 (02/24/2012 14:00 - 02/27/2012 15:36) :

No data received from 2012055+1454 to 2012058+1536 due to server issue

2012060+0230 (retired, lloyd)
trb : 2012046+1900-2012058+1800 (02/15/2012 19:00 - 02/27/2012 18:00) :

Turbidity flat-lining from 2012046+1900 to 2012058+1800

2012058+1413 (retired, lloyd)

No data since 2/24/2012.

2012047+1109 (retired, ussery)

No data received since 2/15/2012 @ 0100.

2012034+1415 (retired, lloyd)
trb : 2011355+1600-2012030+1700 (12/21/2011 16:00 - 01/30/2012 17:00) :

Turbidity data reading high to polling issue

2012017+1122 (retired, ussery)

No data since 2012016+0500. Data has resumed.

2012016+1205 (retired, jrizzo)

Turbidity reading high starting 2011355+1600

2012009+1521 (retired, lloyd)
ssx,vly,pwp,ssy :

Current meter signal strength dropping to 0 from 2011347+1800 to 2012005+2200. Current meter was replaced

2011356+1524 (retired, lloyd)

Backup datasonde was in place while repairs were made to primary datasonde from 2011121+1000 to 2011355+1600. Backup datasonde did not report pH, DO or turbidity.

2011356+1520 (retired, lloyd)

Data transmissions sporadic from 2011347+0200 to 2011349+1418

2011348+1423 (retired, lloyd)

No data from 2011347+0200 to 2011347+1100.

2011333+1153 (retired, ussery)

No data received from current meter since 2011326+1630. Data has resumed.

2011300+1512 (retired, lloyd)

Server did not disseminate data from 2011296+0918 to 2011297+1000

2011272+1915 (retired, ussery)

No data since 2011270+0400. Data has resumed.

2011244+1034 (retired, ussery)

Sporadic data coming in.

2011151+1055 (retired, ussery)

No data from Hydrolab since 5/20/2011.

2011136+1514 (retired, lloyd)

Current meter removed for service on 5/4/2011

2011136+1514 (retired, lloyd)

Datasonde cable in need of replacement

2011117+1102 (retired, ussery)

No data from YSI since 4/19/2011.

2011110+1111 (retired, ussery)

Sonde not reporting turbidity

2011066+1043 (retired, ussery)

Current meter removed for maintenance on 2/14/11. Current meter reinstalled 3/4/2011.

2011046+1917 (retired, ussery)

DO and sat spiking low.

2011046+1649 (retired, lloyd)

No data since 2011045+1900.

2011021+1047 (retired, ussery)

No data since 2011020+0730.

2011019+0427 (retired, lloyd)

No data since 2011018+0500.

2011006+1058 (retired, ussery)

No data since 2011004+0700. Data has resumed.

2011003+2037 (retired, lloyd)

Current meter telemetry suspect starting 20101221

2010356+0501 (retired, lloyd)

No data from sontek since 20101221.

2010337+1440 (retired, lloyd)

DO and sat reading zero.

2010323+1439 (retired, lloyd)

Various YSI reading spiking or flatlining.

2010288+0951 (retired, ussery)

No data since 2010283.

2010127+1143 (retired, ussery)

No data since 2010126+0300.

2010109+1627 (retired, lloyd)

Needs new weight.

2010043+1352 (retired, ussery)

DO flatline @ 0

2009356+1600 (retired, ussery)

DO dropped to 0,

2009303+1244 (retired, ussery)

No data from current meter since 2009301+2236.

2009278+1729 (retired, ussery)

No data received since 2009276+1700.

2009254+1635 (retired, ussery)

No data since 2009254+0412

2008363+1125 (retired, jrizzo)

No data since 2008360. Data resumed.

2008354+2054 (retired, ussery)
spcon sal :

Field calibrated, readings suspect.

2008354+2054 (retired, ussery)

No data from station since 2008352+1800. Service trip planned.

2008352+2159 (retired, lloyd)
2008218+2014 (retired, ussery)

Sporadic data from YSI.

2008193+1256 (retired, ussery)

Sporadic data from YSI. Service trip planned 6/27.

2008151+1343 (retired, ussery)

No data from YSI since 2006146+1500. Data has resumed.

2008070+2025 (retired, ussery)

No data from YSI since 2008067. Service trip planned. No response from YSI, replaced YSI.

2008032+1442 (retired, jrizzo)

Little data over the past 10 hours. Data resumed.

2008029+1517 (retired, ussery)

Current meter % good and wave data suspect. Sensor will be cleaned next service trip.

2008023+1417 (retired, ussery)

Sporadic data received since 2008021+2200. Data is steady.

2008009+1417 (retired, ussery)

Sporadic data since 2008005. Service tripped planned for this week.

2007341+1306 (retired, jrizzo)

No DO. DO temporarily disabled.

2007331+1401 (retired, ussery)

No YSI data since 2007324. Warmup set at 30 seconds data not being downloaded.

2007327+1308 (retired, ussery)

No DO. DO and turbidity have been disable due to faulty readings.

2007313+0427 (retired, ussery)

No data since 2007304+0836. Service trip planned week of Nov. 5th.

2007298+1950 (retired, ussery)

No data since 2007295+1400 but service trip planned Thurs.

2007282+2307 (retired, ussery)

No data since 2007271. Service tripped planned.

2007248+1256 (retired, ussery)

No data from station since 2007238+0900. Service trip planned week of Sept 3.

2007233+1258 (retired, ussery)

No sontek data since 2007225. Sontek and YSI are now being polled via freewave.

2007192+1254 (retired, jrizzo)

X and Y signal strength spiking. Will clean sontek next service trip. Sensor cleaned

2007176+1905 (retired, ussery)

YSI seems to be rebooting, unresponsive. Service trip planned week of June 25th.

2007176+1904 (retired, ussery)

No data from the YSI since 2007168. Service trip planned week of June 25th.

2007169+1255 (retired, ussery)

Sig strength for Vel X spiking to 0, no turb data. Service trip planned for June 11th.

2007135+2138 (retired, jdavis)
ssx sdth swh trb vlx vly pwp ssy sbat :

These series are not coming in because the ysi has been connected directly to the ip modem due to problems with the cube. -- now being polled via separate ip modem named gbra1s

2007101+1115 (retired, jrizzo)

No data since 098. Data resumed.

2007093+1838 (retired, jdavis)

No data since 2007090 - data resumed

2007082+1304 (retired, rperez)

No data since 071. Service trip planned week of March 19.
Station reset, cube and YSI are properly working.

2007039+1151 (retired, jrizzo)

Current meter data incomplete since 035. Data resumed

2007033+1202 (retired, jrizzo)

Battery dropped to 7V. Battery replaced.

2007031+1207 (retired, jrizzo)

No YSI data. Data resumed

2007026+1022 (retired, jrizzo)

No data from the Sontek since 021. Service trip planned week of January 22. Data resumed

2006362+1148 (retired, jrizzo)

Current meter experiencing problems. Problem cleared.

2006352+2121 (retired, jadams)
vlx vly : today (07/18/2018 00:00 - 07/19/2018 00:00) :

meter re-oriented.

2006302+0024 (retired, jadams)
all : 10/27/2006 (10/27/2006 00:00 - 10/28/2006 00:00) :

activated sensor on cube.

2006292+1129 (retired, jrizzo)

Data recovery incomplete. Data improved

2006290+1142 (retired, jrizzo)

No data since 2006284. Problem with YSI or cable. Removed to lab for testing. Sensor reinstalled.

2006276+1020 (retired, jrizzo)

Station struck by lightning, sensors and systems destroyed.

2006276+1020 (retired, jrizzo)

Battery dropped to 9V, Data stopped. Will repair once YSI comes in. YSI replaced.

2006248+1317 (retired, jadams)

Battery not charging

2006234+1457 (retired, duff)

No data since 2006230. Problem was with strange data output from Hydrolab. Modified decoder to deal with it.

2006227+2151 (retired, jdavis)

Sporadic data received since 2006226+1700.

2006215+1412 (retired, kjenkin)

YSI repositioned on platform 6/9/06 until housing is replaced. Housing replaced 2006210

2006215+1411 (retired, kjenkin)

Apparent cross-talk in YSI. YSI replaced with hydrolab

2006211+2359 (retired, kjenkin)

No data since 2006207+0800. Station struck by lightning.

2006191+1224 (retired, jrizzo)

Wave data flat since 2006173. Wave data recovered.

2006161+2301 (retired, kjenkin)

Sontek operation sporadic. Replaced cable

2006100+1054 (retired, jrizzo)

Wave data flatline. Sensor cleaned.

2006053+0047 (retired, kjenkin)

Wave height and period flatline. Sensor cleaned

2005345+1254 (retired, jrizzo)

Turbidity sensor not operaing properly

2005321+1147 (retired, jrizzo)

No data since 2005314.

2005313+1426 (retired, kjenkin)

No data since 2005307+1800. Communication with tower resumed.

2005304+2109 (retired, jrizzo)
bat :

Battery @ 8 and 7v respectively. Battery replaced.

2005296+1122 (retired, jrizzo)

No data since 2005283, Ptocon Tower down. Bad serial port in pc. System repaired

2005285+1112 (retired, jrizzo)

Battery @11V and dropping.

2005285+1112 (retired, jrizzo)

No wave data.

2005277+1125 (retired, jrizzo)

No data since 2005196+1000 from the current meter, waiting on replacement sensor. Sensor replaced.

2005276+1225 (retired, jrizzo)

No data since lighthouse OS upgrade 2005272. Data resumed

2005265+1025 (retired, jrizzo)

No data since 2005253. Polling issue. Tower system replaced.

2005243+1217 (retired, jrizzo)

Sporadic turbidity data since 2005231. Data resumed.

2005207+1451 (retired, kjenkin)

Data recovered daily @ 5:00am and 2:00pm CDT.

2005196+1125 (retired, jrizzo)

Multiple problems began on 2005178 & 2005180. Station serviced, sensor cleaned, problems resolved.

2005164+2047 (retired, kjenkin)

Wave height close to 0. Will service ASAP

2005151+1633 (retired, kjenkin)

Turbidity spiking. Disabled wiper

2005132+2302 (retired, kjenkin)

Velocities flatlining. Cleaned Sontek.

2005130+1325 (retired, kjenkin)

Sontek water depth questionable.

2005101+1036 (retired, jrizzo)

No data since 2005100.

2005088+1944 (retired, kjenkin)

No data since 2005087

2005088+1943 (retired, kjenkin)

Turbidity spiking to 1200. Corrected wiper position

2005061+1626 (retired, duff)

Data sporadic. Problem with computer clock at Port O'Connor tower. Fixed. There should be no gaps from now on.

2005038+2009 (retired, kjenkin)

No data since 2005036. Tower problems

2005032+1418 (retired, kjenkin)

No data since 2005029. Pobot fixed?

2005018+1238 (retired, jrizzo)

Station relocated to new platform 1/11/2005

2005018+1237 (retired, jrizzo)

No data since 2005014. No answer from tower. Tower reset.

2005011+1207 (retired, jrizzo)

No data since 2005008, Port O'Connor Tower not answering. Data recovered.

2005003+1958 (retired, kjenkin)

No 2005 ptocdl folder created - 2005001-003 data lost

2004366+0300 (retired, kjenkin)

No data since 2004361. Files imported from tower.

2004354+0152 (retired, kjenkin)

No data since 2004351. Files retrieved from tower.

2004334+1827 (retired, kjenkin)

Poll not operational at this time. Sonde is self-logging.

2004281+1635 (retired, kjenkin)

Turbidity spiking. Removed spikes. Next service scheduled for 10.08.2004

2004272+0044 (retired, kjenkin)

No data since 2004268. The phone at the Port O'Connor tower doesn't seem to be answering.

2004268+1441 (retired, kjenkin)

Problem processing data to database. Corrected.

2004267+1124 (retired, jrizzo)

No data graphed since 2004265

2004265+1900 (retired, kjenkin)

Problems with Turbidity corrected. Polling command was "confusing" YSI.

2004260+1318 (retired, kjenkin)

New YSI 6600EDS deployed 2004259 (Sept 15, 2004).

2004260+0151 (retired, kjenkin)

YSI deployed

2004223+1918 (retired, kjenkin)

No data since 2004222. Transmissions from tower resumed.

2004216+2128 (retired, kjenkin)
trb : 2004212 (07/30/2004 00:00 - 07/31/2004 00:00) :

YSI turbidity flatlining. Replaced with hydrolab

2004215+1607 (retired, kjenkin)

YSI loaner removed.

2004204+2036 (retired, kjenkin)

DO probe unsteady. Replaced hydrolab

2004189+1934 (retired, kjenkin)

Data not received from tower

2004188+1112 (retired, jrizzo)

No Turbidity data. Turbidity data above 1000NTU, Hydrolab will not collect above 1000NTU.

2004141+1554 (retired, kjenkin)

No data since Day139. Data in log file, trouble connecting to tower.

2004096+1724 (retired, kjenkin)

No data since 096+0530UTC. Data is in log file

2004090+2214 (retired, jrizzo)

Turbidity flatline. Turbidity readings maxed out, sensor may have been fouled, replaced this date, will monitor..

2004083+1432 (retired, ussery)

No data received since day 81. Data is now up to date.

2004078+1756 (retired, jrizzo)

Port O'Connor tower not answering the phone, will visit tower and GBRA1 today.

2004076+1948 (retired, jrizzo)

Battery at 10V. Battery replaced.

2004070+1702 (retired, jrizzo)

No data since 2004069